OpenStudy (anonymous):

Simplify and show work. 4a^2 1 ----- + ----- a^2 a^2

jhonyy9 (jhonyy9):

- so if you want solve it for a , and because the nominator is the same a*2 you can writing 4a*2 +1 = 0 --- 4a*2 = -1 --- a*2 =-1/4 --- because i*2 =(-1) hence (-1)*(1/2) =i --- a_1,_2 =(+,-i)/2 a_1 =i/2 a_2 =(-i)/2